A spinner has 1 through 21.What is the probability that the spinner will stop on an even number or a multiple of 3?
Answer by stanbon(75887)

A spinner has 1 through 21.What is the probability that the spinner will stop on an even number or a multiple of 3?
-----------------------------------
Evens: 2,4,6,8...16,18,20
How many?: 10
--------------------------------
3 multiples: 3,6,9,12,15,18,21
How many?: 7
--------------------------------
Numbers that are both:
6,12,19
How many?: 3
------------------------------
P(even or 3-mult) = P(even) + P(3-mult) - P(even AND 3-mult)
P(even or 3-mult) = 10/21 + 7/21 - 3/21 = 14/21 = 2/3
=================================================================
Cheers,
Stan H.